# KNighter: Transforming Static Analysis with LLM-Synthesized Checkers

### v1

@@ Abstract
Static analysis is a powerful technique for bug detection in critical systems like operating system kernels. However, designing and implementing static analyzers is challenging, time-consuming, and typically limited to predefined bug patterns. While large language models (LLMs) have shown promise for static analysis, directly applying them to scan large codebases remains impractical due to computational constraints and contextual limitations.
@@
We present KNighter, the first approach that unlocks practical LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns. Rather than using LLMs to directly analyze massive codebases, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge. KNighter implements this vision through a multi-stage synthesis pipeline that validates checker correctness against original patches and employs an automated refinement process to iteratively reduce false positives. Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers. To date, KNighter-synthesized checkers have discovered 70 new bugs/vulnerabilities in the Linux kernel, with 56 confirmed and 41 already fixed. 11 of these findings have been assigned CVE numbers. This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
1 Introduction
The reliability of fundamental systems, particularly operating system kernels, depends critically on effective defect detection techniques [23, 38, 12, 7, 21, 28, 35]. Static analysis [4] stands out as a powerful approach that can comprehensively examine source code without execution, enabling the detection of bugs in architecture-specific code, hardware-dependent drivers, and configurations that may be challenging to test in real environments [20, 21, 12, 28]. Unlike dynamic approaches such as fuzzing that require concrete execution environments and can only explore executed paths [2, 38, 9, 36], static analysis can theoretically examine all possible code paths, including those that are difficult or impossible to trigger during runtime. Plus, while formal verification [37, 14, 26] offers mathematical guarantees that static analysis cannot provide, it typically requires even more expertise and manual effort, making it impractical for large-scale systems like operating system kernels.
However, static analysis faces its own set of challenges in practice. Developing effective static analyzers requires deep expertise in both the target system and formal methods, along with substantial engineering effort to implement precise detection mechanisms [12, 7, 1]. This complexity often leads to static analyzers that focus on specific classes of bugs, potentially missing other critical defect patterns.
These observations highlight several key opportunities for advancing static analysis. First, the rich history of the fundamental systems, e.g., the Linux kernel, provides extensive examples of bug patterns that can inform static analysis approaches. Moreover, bug patterns identified in historical patch commits may still exist in the current codebase [12]. Third, beyond manually analyzing and summarizing bug patterns, modern large language models (LLMs) [3, 29, 11] have demonstrated remarkable capabilities in understanding both code and natural language descriptions.
Building on these insights, we propose leveraging LLMs to perform static analysis on real-world systems by learning from historical patch commits. However, directly applying LLMs to analyze real-world systems [38, 15, 18, 42, 10, 33] is challenging due to the enormous size of modern codebases—for example, the Linux kernel comprises over 30 million lines of code. The limited context windows of LLMs make it impractical to process an entire codebase at once, and the associated computational cost is prohibitive. In the case of the Linux kernel, which spans more than 300 million tokens, a single analysis pass using GPT-4o is estimated to cost approximately \$750. Moreover, LLMs are susceptible to hallucination issues, potentially leading to false positives [13, 10, 18]. Besides, LLMs also struggle to understand complex codebases [22].
Our solution addresses these challenges by using LLMs to synthesize static checkers based on historical bug patches rather than directly analyzing source code. These checkers, implemented using Clang Static Analyzer (CSA) [8], can then scan the kernel codebase effectively and efficiently. This approach offers several advantages:
  • It avoids the prohibitive costs and context limitations of processing large codebases directly with LLMs, thereby enhancing scalability.
  • The checkers can naturally evolve alongside kernel development without requiring constant LLM involvement.
  • We can validate the synthesized checkers against the original patches to mitigate hallucination concerns.
  • Our approach provides greater traceability and explainability compared to earlier techniques, as the synthesized checkers directly encode the bug patterns in a transparent, human-readable manner.
However, we observed that current LLMs struggle to synthesize and implement checkers end-to-end—a task that challenges even experienced researchers and developers with expertise in both static analysis and the target codebase. To address this limitation, we develop a multi-stage synthesis pipeline (§ 3.1) that breaks down the complex task into manageable steps. Furthermore, we design a fully automated refinement pipeline (§ 3.2) that iteratively improves the synthesized checkers to reduce false positives, ensuring their practical utility in real-world scenarios.
We implement our approach, KNighter, the first fully automated static analyzer generation pipeline. While our approach is highly generalizable to different software systems, we demonstrate its effectiveness by synthesizing checkers for 61 historical patch commits from the Linux kernel, successfully generating 37 plausible checkers. These checkers demonstrate high quality, with approximately 35% false positive rate. Our evaluation also presents that KNighter-detected vulnerabilities are orthogonal to existing expert-written static analyzers [1]. To date, checkers synthesized by KNighter have detected 70 new bugs/vulnerabilities in the Linux kernel, with 56 confirmed and 41 already fixed. Notably, 11 of these bugs have been assigned CVE numbers.
Our contributions are summarized below:
@@ 1 Introduction
  • New paradigm. We introduce a pioneering LLM-based approach for synthesizing static analyzers from patch commits. To our knowledge, KNighter is the first fully automated static analyzer generation system, establishing a new paradigm for LLM-powered static analysis.
@@
  • End-to-end pipeline. We implement KNighter as a comprehensive system featuring multi-stage synthesis and automated refinement pipelines for the Linux kernel. This design enables detection of diverse bug classes and adapts dynamically to emerging vulnerability patterns.
@@
  • Systematic evaluation. We rigorously evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories. Our results demonstrate that KNighter generates high-precision checkers with practical false positive rates.
@@
  • Real-world impact. KNighter-generated checkers have discovered 70 previously unknown bugs/vulnerabilities in the Linux kernel, with 56 confirmed by developers, 41 already fixed, and 11 assigned CVE numbers—validating both the approach’s effectiveness and practical security value.
Our code is available at ise-uiuc/KNighter.
2 Background and Motivation
@@
(a) Patch for the null-pointer-dereference bug. The pointer returned by devm_kzalloc should be checked.
@@
(c) The checker synthesized by KNighter.
Figure 1: A bug pattern related to devm_kzalloc.
2.1 Clang Static Analyzer
Static analysis [4] inspects code without execution to uncover bugs and vulnerabilities. The Clang Static Analyzer [8] (CSA) employs path-sensitive symbolic execution, constructing an Exploded Graph where each node (an ExplodedNode) captures a ProgramPoint and an abstract ProgramState that maps expressions to symbolic values and memory locations. Its modular design uses checkers—small, event-driven components implemented as subclasses of a Checker template—that register for events (e.g., pre- and post-function calls, dead symbols, pointer escapes) and can extend the ProgramState with custom data via dedicated macros. Developing a new checker involves defining the bug pattern, implementing the appropriate callbacks, registering the checker, and integrating it into the testing framework, all while ensuring clear bug reporting through BugType and BugReport mechanisms.
Take the checker shown in Figure 1(c) as an example. In checkPostCall, the checker first verifies that the function call originates from "devm_kzalloc" using ExprHasName, and if so, it updates the program state by marking the memory region in PossibleNullPtrMap as unchecked. In checkBranchCondition, two patterns are handled: for instance, when a pointer is negated in a condition (e.g., if (!ptr)), the checker marks the region as checked by calling markRegionChecked; similar logic applies for conditions comparing the pointer directly to NULL. The checkLocation callback then looks up the region in PossibleNullPtrMap and, if it remains unchecked, issues a warning through reportUncheckedDereference. Finally, checkBind addresses pointer assignments by updating the PtrAliasMap for both the left-hand side and right-hand side regions, ensuring that any aliasing is properly tracked.
2.2 Motivating Example
@@
Limitations of existing tools. Despite this vulnerability pattern recurring since at least 2021, with our analysis identifying at least four historical patches addressing it, no static analysis tool had been developed to systematically detect these issues. Even specialized kernel checkers like Smatch [1] fail to identify these vulnerabilities because they lack the domain-specific knowledge that devm_kzalloc may return NULL upon failure.
Our approach. KNighter extracts critical insights from the patch: unchecked return values from devm_kzalloc represent potential null-pointer dereference vulnerabilities. The synthesized checker (Figure 1(c)) tracks null-check status across execution paths while correctly handling pointer aliasing—a sophisticated static analysis capability. This checker discovered 3 new vulnerabilities in the Linux kernel. Figure 1(b) presents one such vulnerability exhibiting the same pattern where a null pointer check is missing after a devm_kzalloc call. This bug was subsequently fixed and assigned CVE-2024-50103, confirming the security significance of our approach.
Advantages over direct LLM scanning. Directly using LLMs to scan the Linux kernel would be prohibitively expensive, as devm_kzalloc alone appears over 7K times across 5.4K files. In contrast, KNighter’s static analyzers primarily consume CPU resources rather than repeated LLM invocations, making the approach both scalable and cost-effective. Moreover, since generating the checkers is mostly a one-time effort, they can naturally evolve alongside the codebase.
@@
Figure 2: Overview of KNighter.
3 Design
Terminology. KNighter takes a patch commit as input and outputs a corresponding CSA checker. Valid checkers correctly distinguish between buggy and patched code, flagging pre-patch code as defective while recognizing post-patch code as correct. Plausible checkers [^1] We adopt the term “plausible” from program repair literature [24, 31], where a “plausible” patch passes all test cases and potentially represents the correct fix. are valid checkers that additionally demonstrate practical utility through low false positive rates or a manageable number of reports. We provide formal definitions of these terms in § 4.
Overview. KNighter leverages agentic workflow to process patch commits for static analyzer synthesis, as illustrated in Figure 2. It operates in two phases: checker synthesis (§ 3.1) and checker refinement (§ 3.2). In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using Clang (§ 3.1.3). If compilation errors occur, a syntax-repair agent automatically repairs them based on the error messages. This phase concludes with the generation of valid checkers (§ 3.1.4). In the subsequent checker refinement phase, these valid checkers are deployed to scan the entire codebase for potential bugs. When bug reports are generated, a bug-triage agent evaluates them for false positives, and KNighter refines the checker accordingly. If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and their filtered reports as potential bugs for review.
Figure 3: Patch commit message.
3.1 Checker Synthesis
@@
4       // Iterative checker generation and evaluation
@@
6             // Stage 1: Bug Pattern Analysis
@@
9            // Stage 2: Detection Plan Synthesis
@@
12            // Stage 3: Analyzer Implementation and Repair
@@
15             while hasCompilationErrors(checker) and attempts < maxAttempts do
@@
21                   // Skip evaluation if checker still has errors
@@ 22                   Continue
Algorithm 1 Synthesize checkers with patch commits.
3.1.1 Bug Pattern Analysis
@@
We have developed an LLM-based agent specifically designed to perform this pattern analysis, with the prompt template shown in Figure 4. In addition to the patch, we extract the complete function code that was modified from the kernel codebase. This additional context is crucial because the patch diff alone may not capture all relevant buggy patterns, as some issues depend on the broader context of the code. By providing both the patch and the complete function code to LLMs, we enable a more comprehensive understanding of the bug being patched.
A single bug pattern from a patch can often be expressed and detected in multiple ways, varying in their scope and complexity. Consider the null pointer dereference bug shown in Figure 1(a). One could formulate a general pattern requiring checks after any function that might return a null pointer. While comprehensive, implementing such a broad checker would be challenging even for LLMs. Instead, our approach favors more targeted bug patterns that facilitate precise checker implementation while remaining practical. For the null pointer example, we narrow the scope to specifically verify that pointers returned by devm_kzalloc are checked before dereferencing—a more focused pattern that maintains effectiveness while being significantly more tractable to implement.
3.1.2 Plan Synthesis

### v2

@@ Abstract.
Static analysis is a powerful technique for bug detection in critical systems like operating system kernels. However, designing and implementing static analyzers is challenging, time-consuming, and typically limited to predefined bug patterns. While large language models (LLMs) have shown promise for static analysis, directly applying them to scan large systems remains impractical due to computational constraints and contextual limitations.
@@
We present KNighter, the first approach that unlocks scalable LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns. Rather than using LLMs to directly analyze massive systems, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge. KNighter implements this vision through a multi-stage synthesis pipeline that validates checker correctness against original patches and employs an automated refinement process to iteratively reduce false positives. Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers. To date, KNighter-synthesized checkers have discovered 92 new, critical, long-latent bugs (average 4.3 years) in the Linux kernel; 77 are confirmed, 57 fixed, and 16 have been assigned CVE numbers. This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
1. Introduction
Figure 1. Motivation. Static analysis should scale to address a diverse range of bug patterns and handle massive codebases. Traditional static analysis struggles with covering wide-ranging bugs, whereas LLM-based methods face hurdles in scaling to large codebases.
The reliability of fundamental software systems—particularly operating system (OS) kernels—hinges on robust defect detection methodologies (kernel-fuzzing, ; kernelgpt, ; kernel-static-refcounting, ; kernel-static-cai2023place, ; kernel-static-lyu2022goshawk, ; suzuki2024fitx, ; yang2024whitefox, ; gong2025snowplow, ; gong2023snowcat, ). Among various techniques, static analysis (ayewah2008using, ) stands out for its ability to examine source code without execution, making it indispensable for scenarios involving hardware-dependent drivers, complex or rarely exercised paths, and configurations difficult to reproduce in real environments (kernel-static-lu2019detecting, ; kernel-static-lyu2022goshawk, ; kernel-static-refcounting, ; suzuki2024fitx, ; gong2025snowplow, ; gong2023snowcat, ). Compared to dynamic approaches such as fuzzing—which requires an execution environment and thus only tests actual runtime paths (syzkaller, ; kernelgpt, ; deng2023large, ; yang2023fuzzing, )—static analysis can (in principle) cover all potential execution paths, including corner cases that are seldom triggered in practice. While formal verification techniques (yang2024autoverus, ; lattuada2023verus, ; sun2024clover, ) offer stronger correctness guarantees, their high manual overhead renders them impractical for large-scale systems like OS kernels, making static analysis a more scalable and feasible solution.
The static analysis problem. Large-scale systems present a dual challenge for static analysis: addressing diverse bug patterns and managing enormous codebases, as illustrated in Figure 1. An ideal static analyzer should (i) detect a wide range of defects—including those related to nuanced, system-specific semantics—and (ii) efficiently process millions of lines of code. However, existing techniques typically compromise on one of these critical objectives.
Traditional static analysis. Traditional static analyzers are effective in identifying certain bug types, yet they fundamentally rely on pre-defined, rule-based, or formally modeled checks. This reliance necessitates extensive domain expertise and substantial engineering effort for their development and maintenance (kernel-static-refcounting, ; kernel-static-cai2023place, ; smatch, ). Consequently, these tools are often fine-tuned to a narrow subset of bug patterns, which not only limits their ability to detect unforeseen defects but also hampers their scalability in automatically addressing a broader spectrum of issues.
Emerging LLM-based static analysis. On the other hand, Large Language Models (LLMs) are compelling tools for discovering bug patterns in part because they can learn directly from historical patch commits—a treasure trove of real fixes and associated bug contexts (kernel-static-refcounting, ; seal, ; aphp, ). Their ability to parse both textual and code content (achiam2023gpt, ; team2023gemini, ; grattafiori2024llama, ) suggests that LLMs can adapt to new bug types without explicit rule-crafting. However, directly deploying LLMs on large-scale systems (e.g., the Linux kernel at over 30 million lines of code) confronts severe limitations. Their bounded context windows make it impossible to upload all relevant source code at once, and doing so repeatedly would also incur prohibitive computational costs (potentially hundreds of dollars per thorough scan). In addition, LLMs can hallucinate (huang2025survey, ; du2024vul, ; li2024llm, ), producing plausible but incorrect outputs, especially when faced with the intricacies of large-scale systems (mathai2024kgym, ).
Insight. Can we scale and automate static analysis to handle both diverse bug patterns and enormous codebases? We answer this question by harnessing the strengths of traditional static analysis alongside emerging LLMs. More specifically, we propose synthesizing static checkers using LLMs rather than applying LLM-based analysis directly to the entire codebase. In this paradigm, LLMs learn bug patterns from historical patches, and these insights are encoded into dedicated static analysis checkers. This method circumvents the prohibitive costs and context-length limitations of scanning vast codebases while maintaining the flexibility needed to address a wide spectrum of bugs. Moreover, by validating each synthesized checker against the original patches, we mitigate hallucinations and produce transparent, human-readable logic that developers can trust and maintain.
Technical challenges and our solutions. Although automated checker synthesis holds significant promise, generating complete static analysis logic remains a formidable challenge—even experts struggle with it. To address this, we introduce a multi-stage synthesis pipeline (§ 3.1) that decomposes checker generation into manageable subtasks. Furthermore, to enhance the quality of the synthesized checkers by reducing false positives, we develop a fully automated refinement pipeline (§ 3.2) that leverages bug report triage agents. Together, these pipelines yield checkers that are robust and practical for deployment in real-world scenarios.
We implement our approach in a tool, KNighter, the first fully automated pipeline for synthesizing static analyzers, built upon the open-source Clang Static Analyzer (CSA) (clang-static-analysis, ). Although the methodology generalizes to different systems, our evaluation targets the Linux kernel, one of the most fundamental software systems. From 61 historical patch commits, KNighter successfully synthesizes 37 high-quality checkers with a false positive rate of about 35%. Furthermore, the vulnerabilities detected are orthogonal to those found by existing expert-written analyzers (smatch, ). Demonstrating practical impact, our synthesized checkers have uncovered 92 new, long-latent vulnerabilities (average 4.3 years) in the Linux kernel, resulting in 77 developer confirmations, 57 fixes, and 16 CVEs. These findings validate our approach’s efficacy and its contribution to system reliability.
Our main contributions are summarized as follows:
@@ 1. Introduction
  • Novelty. We introduce a pioneering approach for synthesizing static analyzers from patch commits. To our knowledge, KNighter is the first fully automated static analyzer generation system, establishing a new paradigm for LLM-based static analysis.
@@
  • Approach. We implement KNighter with multi-stage synthesis and automated refinement pipelines for the Linux kernel. This design enables detection of diverse bug classes in large-scale systems.
@@
  • Evaluation. We evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories. Our results demonstrate that KNighter generates high-precision checkers with practical false positive rates.
@@
  • Real-world impact. KNighter-generated checkers have discovered 92 new, long-latent (average 4.3 years) bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 assigned CVE numbers—demonstrating its practical impact on system reliability and security.
KNighter is available at ise-uiuc/KNighter.
2. Background and Motivation
@@
(a) Patch for a Null-Pointer-Dereference bug. The pointer returned by devm_kzalloc should be checked.
@@
(c) A checker synthesized by KNighter for the patch in Fig. 2(a).
Figure 2. A bug pattern related to devm_kzalloc.
2.1. Clang Static Analyzer
Static analysis (ayewah2008using, ) is a technique for detecting bugs by inspecting code without executing it. The Clang Static Analyzer (clang-static-analysis, ) (CSA) serves as a powerful engine for this purpose. It operates using path-sensitive symbolic execution, building an internal representation called an ExplodedGraph. Each node within this graph, an ExplodedNode, represents a specific ProgramPoint paired with an abstract ProgramState. This state meticulously maps program expressions to symbolic values and tracks the contents of memory locations.
The modularity of CSA is built upon checkers. These are small, specialized components, typically implemented as subclasses of a Checker template. Checkers function in an event-driven manner, registering interest in specific analysis events such as pre- and post-function calls, the identification of dead symbols, or instances of pointer escapes. A key capability of checkers is their ability to extend the ProgramState with custom, checker-specific data using provided macros, allowing them to maintain sophisticated state across the analysis. Developing a new checker for CSA generally involves several steps: defining the specific bug pattern to be detected, implementing callback methods corresponding to the relevant analysis events, registering the new checker with the analysis framework, and integrating it into the testing system. Effective bug reporting is crucial, utilizing mechanisms like BugType and BugReport to provide clear diagnostics.
To illustrate, consider the example checker shown in Figure 2(c), which registers four distinct callback functions.
  • The checkPostCall callback activates after function calls. It uses ExprHasName to check if the call was to devm_kzalloc. If so, it updates the custom state map PossibleNullPtrMap to mark the returned memory region as potentially null (unchecked).
  • The checkBranchCondition callback is used to handle conditional checks involving the pointer. It recognizes patterns like negation (if (!ptr)) or direct comparison (if (ptr == NULL)) and updates the state via markRegionChecked to reflect that a null check has occurred for the associated memory region.
  • The checkLocation callback is triggered when a memory location is accessed. More specifically, it consults the PossibleNullPtrMap state; if the region is marked as unchecked at this point, it issues a warning using reportUncheckedDereference.
  • The checkBind callback manages pointer assignments. It updates another custom state map, PtrAliasMap, to track potential aliases between memory regions involved in the assignment, ensuring the checker correctly handles cases where multiple pointers might refer to the same potentially null memory.
2.2. Motivating Example
@@
Limitations of existing tools. Despite this vulnerability pattern recurring since at least 2017 (commit 49af64e), with our analysis identifying at least six historical patches addressing it, no static analysis tool had been developed to systematically detect these issues. Even specialized kernel checkers like Smatch (smatch, ) fail to identify these vulnerabilities because they lack the domain-specific knowledge that devm_kzalloc may return NULL upon failure.
Our approach. KNighter extracts critical insights from the patch: unchecked return values from devm_kzalloc represent potential Null-Pointer-Dereference vulnerabilities. The synthesized checker (written in CSA, Figure 2(c)) tracks null-check status across execution paths while correctly handling pointer aliasing, a sophisticated static analysis capability. This checker discovered 3 new vulnerabilities in the Linux kernel. Figure 2(b) presents one such vulnerability exhibiting the same pattern where a null pointer check is missing for the pointer returned by the devm_kzalloc call. This bug was subsequently fixed and assigned CVE-2024-50103.
Advantages over direct LLM scanning. Directly using LLMs to scan the Linux kernel would be prohibitively expensive, as devm_kzalloc alone appears over 7K times across 5.4K files. In contrast, KNighter’s static analyzers primarily consume CPU resources rather than repeated LLM invocations, making the approach both scalable and cost-effective. Moreover, since generating the checkers is mostly a one-time effort, they can naturally evolve alongside the system.
@@
Figure 3. Overview of KNighter.
3. Design
Terminology. KNighter takes a patch commit as input and outputs a corresponding CSA checker. Valid checkers correctly distinguish between buggy and patched code, flagging pre-patch code as defective while recognizing post-patch code as correct. Plausible checkers [^1] We adopt the term “plausible” from program repair (qi2015analysis, ; xia2023automated, ), where a “plausible” patch passes all test cases and potentially is the correct fix. are valid checkers that additionally demonstrate practical utility through low false positive rates or a manageable number of reports. We provide formal definitions of these terms in § 4.
Overview. KNighter leverages agentic workflow to process patch commits for static analyzer synthesis, as illustrated in Figure 3. It operates in two phases: checker synthesis (§ 3.1) and checker refinement (§ 3.2). In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using CSA (§ 3.1.3). If compilation errors occur, a syntax-repair agent automatically repairs them based on the error messages. This phase concludes with the generation of valid checkers (§ 3.1.4). In the subsequent checker refinement phase, these valid checkers are deployed to scan the entire codebase for potential bugs. When bug reports are generated, a bug-triage agent evaluates them for false positives, and KNighter refines the checker accordingly. If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and the filtered reports as potential bugs for review.
Figure 4. Patch commit message.
3.1. Checker Synthesis
@@
4       # Iterative checker generation and evaluation
@@
6             # Stage 1: Bug Pattern Analysis
@@
9             # Stage 2: Detection Plan Synthesis
@@
12             # Stage 3: Analyzer Implementation and Repair
@@
15             while hasCompilationErrors(checker) and attempts ¡ maxAttempts do
@@
21                   # Skip evaluation if checker still has errors
@@ 22                   Continue
Algorithm 1 Synthesize checkers with input patch.
3.1.1. Bug Pattern Analysis
@@
We have developed an LLM-based agent specifically designed to perform this pattern analysis, with the prompt template shown in Figure 5(a). In addition to the patch, we extract the complete function code that was modified from the kernel codebase. This additional context is crucial because the patch diff alone may not capture all relevant buggy patterns, as some issues depend on the broader context of the code. By providing both the patch and the complete function code to LLMs, we enable a more comprehensive understanding of the bug being patched.
A single bug pattern identified from a patch can be expressed with varying scope and complexity. Consider the Null-Pointer-Dereference involving devm_kzalloc (Figure 2(a)). A broad pattern (e.g., check any potentially null return) is comprehensive, but identifying all relevant functions/conditions poses significant static analysis challenges, hindering robust implementation by LLMs. Consequently, our approach favors more targeted bug patterns derived from the patch context. These facilitate precise and tractable checker synthesis by the LLMs. For the devm_kzalloc example, focusing specifically on its return value yields a targeted pattern that effectively addresses the observed bug class while being significantly more manageable for the LLM to implement correctly compared to the broader, more complex alternative.
3.1.2. Plan Synthesis